The dose makes the poison.

Ethanol is highly toxic if it is consumed in large quantities.

Although it is the second least toxic of the five alcohols discussed in this case, the ingestion of large quantities of ethanol is not safe. One serious problem with ethanol consumption is how the quantity is conveyed to the consumer. For example, American malt liquor can be sold in 23.5 fl. oz.23.5 fl. oz. (695 mL)(695 mL) cans and is 12% volume per volume (v/v) ethanol.

Although these cans are labeled as containing multiple servings, many young people consider one can to be a single serving. Drinking two cans of this beverage can result in an accidental overdose of ethanol.

Two cans may not seem like much until you do the math.

In the previous question, you determined that the total volume of pure ethanol in a 695 mL695 mL can of American malt liquor is 83.4 mL.83.4 mL. To better understand its significance, consider how this volume of ethanol compares with other alcoholic beverages.

Calculate the total volume of ethanol in a standard 25.4 fl. oz25.4 fl. oz (750 mL)(750 mL) bottle of liquor (e.g., whiskey, vodka, rum, etc.), which is typically 40%40% (v/v) ethanol.

Enter your answer in milliliters.

pure ethanol in liquor:

mL

Now calculate the total volume of ethanol in a typical 12 fl. oz12 fl. oz (354.8 mL)(354.8 mL) can of light beer, which is 4% (v/v) ethanol.

Enter your answer in milliliters.

pure ethanol in light beer:

mL
Answer #1

%(vol/vol) =(volume of ethanol/volume of can)×100

Vol of ethanol =( % of ethanol × volume of can)÷ 100

For 750 ml bottle of liquor with 40% (vol/vol) ethanol

Vol of ethanol=(40×750)÷100= 300 ml

Total vol of ethanol in 750 ml bottle with 40% ethanol is 300 ml

For 354.8 ml can of light beer with 4 %(v/v) ethanol

Vol of ethanol= (4×354.8)÷100= 14.192 ml

Total vol of ethanol in 354.8 ml can with 4% ethanol is 14.192 ml

Pure ethanol in liquor= 300 ml

Pure ethanol in light beer= 14.192 ml